OpenStudy (anonymous):

Write an equation in general form for a polynomial function with real coefficients and zeros at 1 and 3 + 5i.

OpenStudy (phi):

to have real coefficients, any complex roots come in complex conjugate pairs. if you have a+ bi then you must also have a - bi (complex conjugate means the imaginary part is negated) so you will need three roots for this problem multiply out (x- a) (x-b) (x-c) where a, b and c are the roots. first step: what are the 3 roots ?

OpenStudy (anonymous):

I,m not 100% sure, but I believe the roots are (x-1) (x+3) (x+5i)

OpenStudy (phi):

x=1 is a root subtract 1 from both sides: x-1= 0 that means (x-1) is one of the factors x= 3 + 5i so (x- (3+5i))=0 or (x-3-5i) =0 and (x-3-5i) is another factor you need the complex conjugate of 3+5i for the 3rd root. It is 3-5i x= 3-5i the 3rd factor is (x-3+5i) now you have to multiply out (x-1)(x-3-5i)(x-3+5i) and remember that i*i = -1